Check p(x) + p(−x) ∈ P
(e)
for every p(x) ∈ R(x). Check that the map
ψ : R[x] → P
(e) given by ψ(p(x)) = p(x)+p(−x)
2
is a linear map. Further, check that
ψ
2 = ψ. Determine the kernel of ψ.
1
Expert's answer
2019-09-25T09:28:24-0400

1)Let "p(x)\in \mathbb R[x]" and "q(x)=p(x)+p(-x)" . We have "q(-x)=p(-x)+p(-(-x))=p(-x)+p(x)=q(x)" , so "q(x)\in P(e)"

2)Check that "\psi" is a linear map. Let "p(x)\in \mathbb R[x]", "q(x)\in\mathbb R[x]" and "\alpha\in\mathbb R", then

"\psi(p(x)+q(x))=\frac{\bigl(p(x)+q(x)\bigr)+\bigl(p(-x)+q(-x)\bigr)}{2}="

"=\frac{p(x)+p(-x)}{2}+\frac{q(x)+q(-x)}{2}=\psi(p(x))+\psi(q(x))" and

"\psi(\alpha p(x))=\frac{\alpha p(x)+\alpha p(-x)}{2}=\alpha \frac{p(x)+p(-x)}{2}=\alpha\psi(p(x))"

So "\psi" is a linear map

3)Check that "\psi^2=\psi". Let "p(x)\in \mathbb R[x]", then

"\psi^2(p(x))=\psi\bigl(\psi(p(x))\bigr)=\frac{\psi(p(x))+\psi(p(-x))}{2}="

"=\frac{\frac{p(x)+p(-x)}{2}+\frac{p(-x)+p(-(-x))}{2}}{2}=\frac{p(x)+p(-x)}{2}=\psi(p(x))"

So "\psi^2=\psi"

4)Let "p(x)\in\ker\psi" , then "0=\psi(p(x))=\frac{p(x)+p(-x)}{2}" . We have "p(x)=-p(-x)", that is "p(x)" is an odd polynomial. So "\ker\psi" is the set of odd polynomials.